TGBT17B_501_05 - Continue Spectral Line Regression Tests, other M&C v18.2 Tests

Goals

  • Test spectral line modes 1, 4, 7, 10, 15, 20, 22, and 29 with position switching
  • Test all spectral line modes with frequency switching except 1, 4, 7, and 10 (already got those as part of TGBT17B_501_04)
  • Test GUPPI c0800x0128 and c0800x0256 stand alone on B1937+21 in search mode to see if broad pulse shows up again
  • Same as above with VPM and GUPPI
  • Do an AutoOOF sequence
  • If time, test i0800x0512 at C-Band

Details

  • Paul M starts switch to v18.2 at 15:00
  • Switch over finishes at 15:13
  • Starting off with frequency switched scans. Will do an AutoPeakFocus to test some GFM changes Paul M made
  • Using AcceptanceTestsSpecLineFSW SB
    • SB submitted at 15:20
    • Scans #1--5 are peak/focus
      • There was a second peak and some sharp features in the pointing and scans. Trying a second AutoPeakFocus on a different source.
    • Scans #6-10 are peak/focus
      • Still seeing some sharp spikes but no second broad source signature. Maybe RFI? Paul analyzed the scans using a 16.3 version of GFM and got the same results.
    • Continuing on with planned FSW observations
    • Scan #11 is Mode 2 FSW
      • Looks good
    • Scan #12 is Mode 3 FSW
      • Looks good
    • Scan #13 is Mode 5 FSW
      • Looks good
    • Scan #14 is Mode 6 FSW
      • Looks good
    • Scan #15 is Mode 6 FSW (switched to 60 s scans)
      • Looks good
    • Scan #16 is Mode 8 FSW
      • Looks good
    • Scan #17 is Mode 9 FSW
      • Looks good
    • Scan #18 is Mode 11 FSW
      • Looks good
    • Scan #19 is Mode 12 FSW
      • Looks good
    • Scan #20 is Mode 13 FSW
      • Looks good
    • Scan #21 is Mode 14 FSW
      • Looks good
    • Scan #22 is Mode 15 FSW
      • Looks good
    • Scan #23 is Mode 16 FSW
      • Looks good
    • Scan #24 is Mode 17 FSW
      • Looks good
    • Scan #25 is Mode 18 FSW
      • Looks good
    • Scan #26 is Mode 19 FSW
      • Looks good
    • Scan #27 is Mode 20 FSW
      • Looks good
    • Scan #28 is Mode 21 FSW
      • Looks good
    • Scan #29 is Mode 22 FSW
      • Looks good
    • Scan #30 is Mode 23 FSW
      • Looks good
    • Scan #31 is Mode 24 FSW
      • Looks good
    • Scan #32 is Mode 25 FSW
      • Looks good
    • Scan #33 is Mode 26 FSW
      • Looks good
    • Scan #34 is Mode 27 FSW
      • Looks good
    • Scan #35 is Mode 28 FSW
      • Looks good
    • Scan #36 is Mode 29 FSW
      • Looks good
  • Moving on to position switched scans using AcceptanceTestsSpecLineTP
    • SB submitted at 16:35
    • Scans #37/38 are Mode 1 on/off pair
      • Looks good
    • Scans #39/40 are Mode 4 on/off pair
      • Looks good
    • Scans #41/42 are Mode 7 on/off pair
      • Looks good
    • Scans #43/44 are Mode 10 on/off pair
      • Looks good
    • Scans #45/46 are Mode 15 on/off pair
      • Looks good
    • Scans #47/48 are Mode 20 on/off pair
      • Looks good
    • Scans #49/50 are Mode 22 on/off pair
      • Looks good
    • Scans #51/52 are Mode 29 on/off pair
      • Looks good
  • Now testing GUPPI c0800 search modes on B1937+21
    • SB submitted at 17:01
    • Scan #53 is c0800x0128 search mode
      • Pulsar not detected
    • Scan #54 is c0800x0256 search mode
      • Pulsar not detected
    • Scan #55 is c0800x0512
      • Pulsar there but broad
      • No error in IF setup that I can see
  • B1937+21 set. Going to move to J0218+4232
    • Scan #56 is c0800x0128 search
      • Pulsar not very bright. Tough to tell what's happening.
    • Scan #57 and #58 are c0800x0256 and c0800x0512 search
      • Again, not able to tell much
  • Going to try J1125+7819
    • Scan #59 is c0800x0128 search
      • No obvious issues compared with earlier NANOGrav scan
    • Scan #60 is c0800x0256 search
      • No obvious issues compared with earlier NANOGrav scan
    • Scan #61 is c0800x0512 search
      • No obvious issues compared with earlier NANOGrav scan
  • Taking c0800 data with VPM and GUPPI
    • Scans #63--65 are c0800x0128, 0256, and 0512
      • GUPPI scans look OK. Some of the VPM scans had levels that were too low
  • Testing C-Band i0800 modes
    • Scans #66-68 are i0800x0512 cal, fold, and search
      • No ripples
    • Scans #69-71 are i0800x1024 cal, fold, and search
      • No ripples
    • Scans #72-74 are i0800x2048 cal, fold, and search
  • So it turns out the noise cal was in the previous 1500 MHz VPM scans. Taking one scan each in i1500x0512 and c1500x0512 w/out the cal to see if that gets rid of ripples
    • Scans #75--77 are i1500x0512 cal, fold, and search
      • No ripples
    • Scans #78--80 are c1500x0512 cal, fold, and search
      • Didn't see pulsar in fold scan. Need to look more closely tomorrow
  • No going to try a Q-Band AutoOOF.
    • Tried to configure using the tp_config_multi_beam setup from /home/astro-util/projects/GBog/configs/tp_config_mutli_beam.py but LO1 was resetting to the previous rest frequency and then throwing an Abort.
    • Continued anyway and when it configured for an AutoPeakFocus the LO issue cleared
    • Weather is definitely not good for Q-Band. Pointing scans look awful. But will forge ahead.
    • OOF processing failed. Second round of peak scans actually don't look too terrible
      • Paul found a sparrow bug that caused the OOF to fail.
  • Going to try an Argus AutoPeakFocus and come back to OOF if time.
    • Scan aborted. Astrid reported "Error processing scan 94 : , database "DCR" not define in current context". Also got an Error in Argus: "The cold IF state is OFF. Try running the startup script again. Aborting scan." Also "Bias current for cold IF amplifiers is out of range."
  • Paul fixed some bugs, so going to try the OOF again at Q-Band.
    • This time it seems to have worked. RMS is 1162 microns
    • Peak scans look pretty good this time
  • Paul wants to try Argus one more time
    • Same errors
  • Paul begins switching back at 19:58
  • Mode 29 checkout was OK. Antenna didn't go into running but the VEGAS scan seemed fine. Will monitor next project startup but I am considering this done
  • Sessions ends at 20:30

Conclusions

  • Overall seemed successful. Will look at data this week.
-- RyanLynch - 2018-02-11
Topic revision: r4 - 2018-02-11, RyanLynch
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback